[Hallc_running] [New Logentry] RC Daily Update (Oct 04, 2021)

cyero at jlab.org cyero at jlab.org
Mon Oct 4 07:50:03 EDT 2021

Logentry Text:
  ************************************************************* <br>
  <mark> ANNOUNCEMENT: Upcoming Target Training Opportunities </mark> <br>
  ************************************************************* <br>

  <li> <u>On behalf of Silviu Covrig Dusa: </u>  <br>
    There will be an online LH2 target training on Monday Oct 4 starting at 1pm. <br>
    The training should not last more than 90 min and will be conducted via bluejeans. <br>
    This training is required of people who'd like to take target operator shifts and who have <br>
    not yet been LH2 target trained in the 12 GeV era. <br>
    The link to the bj meeting is below. <br>
    <a href="https://bluejeans.com/127830759/1614?src=calendarLink"> https://bluejeans.com/127830759/1614?src=calendarLink </a>

<b>Accelerator Status:</b>
  <li> Yesterday, Accelerator South Linac "2K Colbox Trip" for a second time caused loss of beam for a few hours (~11:00 - 05:00) yesterday, but beam quickly recovered and has been very stable. </li>
    <li> See <a href="https://logbooks.jlab.org/entry/3917032">Oct 03 Day Beam Summary</a>,  <a href="https://logbooks.jlab.org/entry/3917211">Oct 03 Swing Beam Summary</a>, <a href="https://logbooks.jlab.org/entry/3917461">Oct 04 Owl Beam Summary</a>  </li>

<b>General Hall C Status:</b>
  <li> Yesterday early in the morning (~9:00-10:00), SHMS was rotated to 5.505 deg (a milestone for this experiment and for the Hall). </li>
  <li> During the "Controlled Access" to rotate spectrometer, Brad installed a temporary replacement camera for HCCAM03 which had failed a few days ago --> <a href="https://logbooks.jlab.org/entry/3915871">HCCAM03 Broken</a>, <a href="https://logbooks.jlab.org/entry/3916985">HCCAM03 replaced</a></li>
  <li> No showstoppers reported since beam recovered from the "CHL 2K Cold Box Trip". </li>

<b>Hall C PION-LT Data Collection/Analysis Status:</b><br>
  <li> As of 05:50 today (Owl Shift), we have accumulated: 791 Pions (7.91 %) and 1469.491 mC since SHMS was rotated. </li>

We will have brief RC Daily Meeting today to discuss plan on moving forward.

Bluejeans remote connection information:  <br>
To join the Meeting: <br>
 <a href="https://bluejeans.com/861439920">https://bluejeans.com/861439920</a> <br>
To join via Room System: <br>
Video Conferencing System: bjn.vc -or- <br>
Meeting ID : 861439920 <br>
To join via phone : <br>
1) Dial: +1.408.740.7256 (US (San Jose)) <br>
+1.888.240.2560 (US Toll Free)) <br>
+1.408.317.9253 (US (Primary, San Jose)) <br>
(see all numbers -  <a href="http://bluejeans.com/numbers">http://bluejeans.com/numbers</a> ) <br>
2) Enter Conference ID : 861439920 <br>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3917492
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211004/f891af8f/attachment.html>

More information about the Hallc_running mailing list